# “In Dialogues We Learn”: Towards Personalized Dialogue Without Pre-defined Profiles through In-Dialogue Learning

### v2

@@ Abstract
Personalized dialogue systems have gained significant attention in recent years for their ability to generate responses in alignment with different personas. However, most existing approaches rely on pre-defined personal profiles, which are not only time-consuming and labor-intensive to create but also lack flexibility. We propose In-Dialogue Learning (IDL), a fine-tuning framework that enhances the ability of pre-trained large language models to leverage dialogue history to characterize persona for completing personalized dialogue generation tasks without pre-defined profiles. Our experiments on three datasets demonstrate that IDL brings substantial improvements, with BLEU and ROUGE scores increasing by up to $200\%$ and $247\%$, respectively. Additionally, the results of human evaluations further validate the efficacy of our proposed method.
@@ 1 Introduction
Recently, there has been growing interest in building personalized dialogue systems Tang et al. (2023); Chen et al. (2023c); Huang et al. (2023); Chen et al. (2023a); Tu et al. (2022). Such systems are often adept at incorporating special personal characteristics into responses. Consequently, personalized dialogue systems offer enhanced flexibility, enabling them adapt more effectively to a wide range of conversational scenarios, such as role-playing games Park et al. (2023).
@@
To achieve personalized dialogues, a common practice is to condition a dialogue model on a profile that explicitly depicts the personality traits one aims to portray with a textual description Song et al. (2021); Liu et al. (2022); Chen et al. (2023b). While a profile can effectively delineate the desired personality traits, creating an accurate profile is nevertheless time-consuming and arduous.
In this work, we attempt to develop a model capable of performing personalized dialogue generation without the need of profiles designed in advance. To this end, we introduce In-Dialogue Learning (IDL), a two-stage framework that directly learns persona information from dialogue sessions, and leverages the learnt insights to synthesize responses that exhibit explicit personality characteristics (cf., Figure 1).
@@
IDL comprises a Mutual Supervised Learning (MSL) stage and a Deep Personalized Alignment (DPA) stage. The objective of MSL is to equip a dialogue model with persona knowledge conveyed in dialogue sessions. To this end, one can simply select one dialogue as the target and take the remaining as the reference to perform few-shot learning to optimize the dialogue model. Such a straightforward implementation, however, suffers from two major problems: (1) unified reference dialogues normally contain abundant irrelevant information to the target dialogue, which increases the difficulty of learning; and (2) incoherent transition in multiple dialogues could cause disruption in the dialogue structure. To address the problems, we propose Static Persona Identification (SPI) and Dynamic Persona Identification (DPI) to cluster and re-order dyadic dialogues between a target person and the other interlocutors for effective IDL. SPI divides the dialogues of the person into multiple persona-relevant clusters, ensuring that the target dialogue can easily access inter-session personalized information from reference dialogues from each cluster. DPI further re-orders the reference dialogues by minimizing the gaps in these dialogues, which is measured by conversational edit distance (convED) Lavi et al. (2021).
@@
To enhance the alignment of responses with the target persona Ouyang et al. (2022); Yuan et al. (2023); Song et al. (2023); Hong et al. (2023), we adopt reinforcement learning through the Deep Personalized Alignment stage. We introduce Direct Preference Optimization with Criterion (DPOC), an optimization method derived from DPO Rafailov et al. (2023) to mitigate preference degradation problem with a criterion-based penalty. This approach ensures that responses are more closely aligned with the target persona learned from reference dialogues.
@@
We conducted experiments on several personalized dialogue datasets to evaluate the effectiveness of IDL. Evaluation results show that IDL achieves performance comparable to very strong profile-based methods, without utilizing any pre-defined profile information and supervision. In comparison to traditional personalized dialogue approaches, IDL demonstrates significant improvements, highlighting the benefits of leveraging large language models for personalized dialogue. Furthermore, IDL shows significant improvement over ICL when both utilize large language models, with BLEU and ROUGE scores increasing up to $200\%$ and $247\%$, respectively. This suggests that, unlike ICL, which primarily learns from data samples, IDL is more effective at incorporating persona information within dialogues.
@@ 3 Method
We present technique details of In-Dialogue Learning (IDL) in this section. As shown in Figure 2, IDL involves two stages: Mutual Supervised Learning (MSL) and Deep Personalized Alignment (DPA). In the MSL stage, we propose static and dynamic persona identification to cluster and re-order the dialogues of the target person, and then organize these dialogues into an end-to-end form to perform supervised learning, endowing the model with the ability to leverage persona information within previous dialogues. In the DPA stage, we further extend the DPO algorithm with Criterion (abbreviated as DPOC) to address the issue of preference degradation through the incorporation of criterion examples and penalty terms, facilitating fine-grained personalized learning.
@@ 3.1 Problem Formalization
The goal of IDL is to generate responses that reflect the personality of a target person $u$ based on his/her previous dialogues $\mathbb{D}^{u}$. Formally, $\forall d^{(u,v)}=(q_{1},r_{1},\ldots,q_{t},r_{t})\in\mathbb{D}^{u}$, $d^{(u,v)}$ represents a dialogue between $u$ and another participant $v$ where $(q_{i},r_{i})$ is the $i$-th turn with $q_{i}$ the utterance from $v$ and $r_{i}$ the response from $u$, respectively. Given the current dialogue context $C_{i}=(q_{1},r_{1},\dots,q_{i})$, the generation of IDL can be formulated as
@@
where LM represents the language model, and $\Theta$ is the learnable parameters. Following the common practice, we concatenate $\mathbb{D}^{u}$ and $C_{i}$ as the input of the LM.
@@ 3.2.1 Static Persona Identification
We employ a public dataset PersonaExt Zhu et al. (2023) and train a persona extractor to recognize persona-intensive utterances in a dialogue corpus. PersonaExt segregates persona information within dialogues into triples of <subject, relationship, object>. The dataset defines $105$ types of relationships. Based on the dataset, we develop the persona extractor (abbreviated as Ext) to directly extract the triples from the dialogue. Then, the extracted objects are used to locate the persona-intensive utterances. We formulate the extraction process as
$\{p_{j}^{(u,v)}\}_{j=1}^{n}=\text{Ext}(d^{u,v}),$ (2)
where $p_{j}^{(u,v)}$ denotes a persona-intensive utterance in dialogue $d^{u,v}$. The extracted utterances are then transformed to a vector $z^{(u,v)}$ by
@@
where we utilize the sentence-embedding model as the Enc [^1] https://huggingface.co/sentence-transformers/all-mpnet-base-v2. Based on $\{z^{(u,v)}\}$ and the euclidean metric, $\mathbb{D}^{u}$ is clustered by k-means algorithm:
$K^{u}=\text{KMeans}(\{z^{(u,i)}\},c),$ (4)
where $c$ is the number of clusters. Subsequently, within each cluster $K_{j}^{u}\in K^{u},j=1,2,\dots,c$, we randomly select a dialogue as the target dialogue while the closest top-$k$ in the remaining dialogues are regarded as the reference dialogues.
@@ 3.2.2 Dynamic Persona Identification
To quantify the semantic distance between dialogues, we introduce Conversation Edit Distance (convED) Lavi et al. (2021). The convED metric is akin to the traditional edit distance, but it modifies the basic unit of editing from characters to sentences within a dialogue. The metric aligns one dialogue with another through the processes of inserting, deleting, and substituting sentences. Detailed formulations of convED are presented in Appendix A.2.
@@
In each cluster of $K^{u}$, we concatenate the reference dialogues according to the optimal order and split the target dialogue with the last utterance as a response and the remaining as the context. These data elements satisfy Equation 1, and we can optimize the LM by minimizing the negative likelihood loss. Above processes endow the model with basic IDL ability, which could generate personalized responses based on reference/historical dialogues.
@@ 3.3 Deep Personalized Alignment
The model after MSL initially exhibits the ability of personalized response generation by referencing some dialogues. However, due to hallucinations of LLMs Kalai and Vempala (2023) and complexity of long context, it still fall short in generating personalized response in a more precise manner. Consequently, we introduce the preference alignment technique into IDL for Deep Personalized Alignment (DPA).
@@ 3.3.1 DPOC
The first consideration for preference alignment is Direct Preference Optimization (DPO) Rafailov et al. (2023). DPO distinguishes itself from conventional reinforcement learning algorithms by bypassing the need for reward models, thereby conserving time and computational resources and reducing the complexity typically associated with reinforcement learning. However, DPO encounters a challenge in the form of unstable training outcomes. This instability arises because the primary objective of DPO is to widen the gap between chosen and rejected examples, while it overlooks the diminishing rewards of the chosen examples. Thus, even when the disparity between chosen and rejected examples increases, it may be caused by a concurrent decrease in rewards for both chosen and rejected examples, ultimately leading to a diminished efficacy of the optimized model. This issue is referred as preference degradation.
@@
The criterion sample reward $r_{crt}$ typically serve as intermediary benchmarks between chosen sample reward $r_{cho}$ and rejected sample reward $r_{rej}$. They offer a reference point for the optimization process in DPOC. Specifically, if the reward from a chosen sample falls below that of a criterion sample, or if the reward of a rejected sample’s reward is unexpectedly high compared to criterion examples, the current model incurs a penalty, which is represented by $\mathcal{P}(r_{cho},r_{crt})$ and $\mathcal{P}(r_{crt},r_{rej})$, respectively. This mechanism contributes to alleviating the preference degradation problem.
@@ 3.3.2 Data Construction
In the context of personalized dialogue, we identify three distinct types of criterion examples (cf., Figure 2 right). Each of them utilizes persona information with inaccuracies. (1) Inconsistency: includes information conflicting with the persona established in the dialogue sessions. (2) Fabrication: introduces personality details not mentioned in the dialogue sessions. (3) Inversion: adopts the persona information of the other participant. Given dialogue sessions $\mathbb{D}^{u}$, the context of on-going dialogue context $C$ and a chosen sample $h_{cho}$ of the current response, the construction of the three types of criterion examples are detailed as follows:
[…]
Fabrication. We encode sentences in the dataset, selecting top-$m$ candidates with highest semantic similarity to $h_{cho}$. A candidate, $h_{crt}$, is randomly chosen ensuring $\text{Ext}(h_{crt})\cap\text{Ext}(\mathbb{D}^{u})=\emptyset$. For example, from the utterance “My hobbies are watching movies and riding bicycles”, we extract triples <I, hobby, watching movies> and <I, hobby, riding bicycles>. As the triples are not involved in $\text{Ext}(D^{u})$, we can adopt this utterance as $h_{crt}$.
@@ 4.1 Datasets
ConvAI2 Dinan et al. (2020) is a high-quality English dataset focused on personalized dialogues. Each dialogue revolves around a specific profile. The dataset is expanded from the classic PersonaChat Zhang et al. (2018) by crowd workers.
Cornell Movie-Dialogs Corpus Danescu-Niculescu-Mizil and Lee (2011) contains over $220,000$ dialogues collected from more than $600$ movies with rich meta-data, offering a diverse range of dialogues between $10,000$ pairs of characters.
LIGHT Urbanek et al. (2019) is a large-scale crowdsourced fantasy text adventure game research platform. We extract dialogues of each character to form the dataset used in the experiments.
@@ 4.2 Baselines
Large Language Models have made great progress in recent years. We select LLaMA-2-7B-Chat and LLaMA-2-13B-Chat Touvron et al. (2023) as the backbones of IDL, and name the models LLaMA-2-7B IDL and LLaMA-2-13B IDL, respectively. Besides, Vicuna [^2] https://lmsys.org/blog/2023-03-30-vicuna/ and WizardLM Xu et al. (2023) are involved in comparison, where the former is an open-source chatbot developed by fine-tuning LLaMA with user-shared conversations sourced from ShareGPT, and the latter is fine-tuned from LLaMA-2, starting with a basic set of instructions.
Since profiles are available in ConvAI2, we compare IDL with the profile-based approaches as well as the the profile-free approaches on this dataset. As existing profile-based approaches are not based on LLMs, we further fine-tune LLaMA-2-7B-Chat and LLaMA-2-13B-Chat with the gold profiles in ConvAI2 for fair comparison, and name the models LLaMA-2-7B gold and LLaMA-2-13B gold, respectively. On Movie and LIGHT, we assess the transferability of IDL by comparing LLaMA-2-7B IDL and LLaMA-2-13B IDL, both fine-tuned on ConvAI2, against other LLMs utilizing in-context learning method.

### v4

@@ Abstract
Personalized dialogue systems have gained significant attention in recent years for their ability to generate responses in alignment with different personas. However, most existing approaches rely on pre-defined personal profiles, which are not only time-consuming and labor-intensive to create but also lack flexibility. We propose In-Dialogue Learning (IDL), a fine-tuning framework that enhances the ability of pre-trained large language models to leverage dialogue history to characterize persona for personalized dialogue generation tasks without pre-defined profiles. Our experiments on three datasets demonstrate that IDL brings substantial improvements, with BLEU and ROUGE scores increasing by up to $200\%$ and $247\%$, respectively. Additionally, the results of human evaluations further validate the efficacy of our proposed method.
@@ 1 Introduction
Recently, there has been growing interest in personalized dialogue systems Tang et al. (2023); Chen et al. (2023c); Huang et al. (2023); Chen et al. (2023a); Tu et al. (2022). Such systems are often adept at incorporating special personal characteristics into responses. Consequently, they offer enhanced flexibility, enabling them to adapt more effectively to a wide range of conversational scenarios, such as personal assistants or chatbots  [^1] https://character.ai/.
@@
A common practice in personalized dialogues is to condition a dialogue model on a pre-defined profile that explicitly depicts the personality traits one aims to portray with a textual description. While there have been extensive studies along this line Song et al. (2021); Liu et al. (2022); Chen et al. (2023b), we explore the problem from a different angle: instead of using a brief profile to describe a person’s personality, we leverage multiple conversations between the individual and others to build a personalized dialogue model. Consequently, the model can generate personalized dialogues without the need for pre-designed profiles which could be both time-consuming and labor-intensive. Furthermore, as dialogue history accumulates, past conversations may provide more personalized information than a static profile.
@@
We introduce In-Dialogue Learning (IDL), a two-stage framework that directly learns persona information from dialogue sessions, and leverages the learnt insights to synthesize responses that exhibit explicit personality characteristics (cf., Figure 1). IDL comprises a Mutual Supervised Learning (MSL) stage and a Deep Personalized Alignment (DPA) stage. The objective of MSL is to equip a dialogue model with persona knowledge conveyed in dialogue sessions. To this end, one can simply select one dialogue as the target and take the remaining as the reference to perform few-shot learning to optimize the dialogue model. Such a straightforward implementation, however, suffers from two major problems: (1) unified reference dialogues normally contain abundant irrelevant information to the target dialogue, which increases the difficulty of learning; and (2) incoherent transition in multiple dialogues could cause disruption in the dialogue structure. To address the problems, we propose Static Persona Identification (SPI) and Dynamic Persona Identification (DPI) to cluster and re-order dyadic dialogues between a target person and the other interlocutors for effective IDL. SPI divides the dialogues of the person into multiple persona-relevant clusters, ensuring that the target dialogue can easily access inter-session personalized information from reference dialogues from each cluster. DPI further re-orders the reference dialogues by minimizing the gaps in these dialogues, which is measured by conversational edit distance (convED) Lavi et al. (2021).
@@
To better align responses with the target persona Ouyang et al. (2022); Yuan et al. (2023); Song et al. (2023); Hong et al. (2023), we introduce Direct Preference Optimization with Criterion (DPOC), an optimization method derived from DPO Rafailov et al. (2023) to mitigate preference degradation problem with a criterion-based penalty. This approach ensures that responses are more closely aligned with the target persona learned from reference dialogues.
@@
We conduct experiments on several personalized dialogue datasets to evaluate the effectiveness of IDL. Evaluation results show that IDL achieves performance comparable to very strong profile-based methods, without utilizing any pre-defined profile information. In comparison to traditional personalized dialogue approaches, IDL demonstrates significant improvements, highlighting the benefits of leveraging large language models for personalized dialogue. Furthermore, IDL shows significant improvement over In-Context Learning (ICL) when both utilize large language models, with BLEU and ROUGE scores increasing up to $200\%$ and $247\%$, respectively. This suggests that, unlike ICL, which primarily learns from data samples (single-turn), IDL is more effective at incorporating persona information within dialogues (multi-turn).
@@ 3 Method
As shown in Figure 2, In-Dialogue Learning (IDL) involves two stages: Mutual Supervised Learning (MSL) and Deep Personalized Alignment (DPA). In the MSL stage, we propose static and dynamic persona identification to cluster and re-order the dialogues of the target person, and then organize these dialogues into an end-to-end form to perform supervised learning, endowing the model with the ability to leverage persona information within previous dialogues. In the DPA stage, we further extend the DPO algorithm with Criterion (abbreviated as DPOC) to address the issue of preference degradation through the incorporation of criterion examples and penalty terms, facilitating fine-grained personalized learning.
@@ 3.1 Problem Formalization
The goal of IDL is to generate responses that reflect the personality of a target person $u$ based on his/her previous dialogues $\mathbb{D}^{u}$. Formally, $\forall d^{(u,v)}=(q_{1},r_{1},\ldots,q_{t},r_{t})\in\mathbb{D}^{u}$, $d^{(u,v)}$ represents a dialogue between user $u$ and the other participant $v$ where $(q_{i},r_{i})$ is the $i$-th turn with $q_{i}$ the utterance from $v$ and $r_{i}$ the response from $u$, respectively. Given the current dialogue context $C_{i}=(q_{1},r_{1},\dots,q_{i})$, the generation of IDL can be formulated as
@@
where LM represents the language model, and $\Theta$ is the learnable parameters. Following the common practice, we concatenate $\mathbb{D}^{u}$ and $C_{i}$ as the input.
@@ 3.2.1 Static Persona Identification
We employ a public dataset PersonaExt Zhu et al. (2023) and train a persona extractor to recognize persona-intensive utterances in a dialogue corpus. PersonaExt segregates persona information within dialogues into triples of <subject, relationship, object>. The dataset defines $105$ types of relationships. Based on the dataset, we develop the persona extractor (abbreviated as Ext) that can directly extract the above-mentioned triples from a dialogue by fine-tuning an LLM. More details about training of the persona extractor are presented in Appendix A.2. The persona extractor identifies and extracts persona-intensive utterances from a dialogue by recognizing utterances that contain at least one object in the extracted triples. Formally, the extraction process can be formulated as
$\{p_{j}^{(u,v)}\}_{j=1}^{n}=\text{Ext}(d^{(u,v)}),$ (2)
where $p_{j}^{(u,v)}$ is a persona-intensive utterance in dialogue $d^{(u,v)}$, with $u$ and $v$ as the participants. $\{p_{j}^{(u,v)}\}_{j=1}^{n}$ are then encoded as the persona representation $z^{(u,v)}$ of $d^{(u,v)}$ by
@@
where Enc()($\cdot$)( ⋅ ) is an off-the-shelf sentence encoder [^2] https://huggingface.co/sentence-transformers/all-mpnet-base-v2. Based on $\{z^{(u,v)}\}$, dialogues $\{d^{(u,v)}|d^{(u,v)}\in\mathbb{D}^{u}\}$ are clustered by k-means algorithm [^3] We tested various clustering algorithms, including k-means, BSCAN, Mean Shift, WARD, and BIRCH, but observed no significant differences. Therefore, we chose k-means for its simplicity and widespread use.:
$K^{u}=\text{KMeans}(\{z^{(u,v)}\},c),$ (4)
where $z^{(u,v)}$ serves as the index of dialogue $d^{(u,v)}$ and $c$ is the number of clusters. Subsequently, within each cluster $K_{j}^{u}\in K^{u},j=1,2,\dots,c$, we randomly select a dialogue as the target dialogue while the closest top-$k$ in the remaining dialogues are regarded as the reference dialogues.
@@ 3.2.2 Dynamic Persona Identification
To quantify the semantic distance between dialogues, we introduce Conversation Edit Distance (convED) Lavi et al. (2021). It is akin to the traditional edit distance, but it modifies the basic unit of editing from characters to sentences within a dialogue. The metric aligns one dialogue with another through the processes of inserting, deleting, and substituting sentences. Detailed formulations of convED are presented in Appendix A.3.
@@
In each cluster of $K^{u}$, we concatenate the reference dialogues according to the optimal order and split the target dialogue with the last utterance as a response and the remaining as the context. These data elements satisfy Equation 1, and we can optimize the LM by minimizing the negative likelihood loss. The above processes endow the model with basic IDL ability, which could generate personalized responses based on reference dialogues.
@@ 3.3 Deep Personalized Alignment
The model after MSL initially exhibits the ability of personalized response generation by referencing some dialogues. However, due to hallucinations of LLMs Kalai and Vempala (2023), it still falls short in generating more precise personalized responses. Consequently, we propose Deep Personalized Alignment (DPA) in IDL.
@@ 3.3.1 DPOC
Previous DPO Rafailov et al. (2023) method for preference alignment encounters a challenge in the form of unstable training outcomes. This instability is aroused by the primary objective of DPO, which maximizes the generation probability gap between chosen and rejected examples. This objective will overlook the diminishing rewards of the chosen examples. Thus, even when the disparity between chosen and rejected examples increases, it may be caused by a concurrent decrease in rewards for both chosen and rejected examples, ultimately leading to a diminished efficacy of the optimized model. This issue is referred as preference degradation.
@@
The criterion sample reward $r_{crt}$ typically serves as intermediary pivots between chosen sample reward $r_{cho}$ and rejected sample reward $r_{rej}$. Specifically, if the reward from a chosen sample falls below that of a criterion sample, or if the reward of a rejected sample’s reward is unexpectedly high compared to criterion examples, the current model incurs a penalty, which is represented by $\mathcal{P}(r_{cho},r_{crt})$ and $\mathcal{P}(r_{crt},r_{rej})$, respectively. Detailed formulations are presented in Appendix A.4.
@@ 3.3.2 Data Construction
To perform DPOC, we need to specify the criterion samples. The intuition of criterion sample construction comes from analysis of the model after the MSL stage, where we observe three major problems, including responses revealing fictitious persona information, conflicting with the persona set by the context, and confusing the partner’s persona with the target person’s. Based on the analysis, we consider the following three types of criterion samples (cf., Figure 2 right): (1) Inconsistency: includes information conflicting with the persona established in the dialogue sessions. (2) Fabrication: introduces personality details not mentioned in the dialogue sessions. (3) Inversion: adopts the persona information of the other participant.
Given dialogue sessions $\mathbb{D}^{u}$, the context of ongoing dialogue $C$ and a chosen sample $h_{cho}$ of the current response, the construction of the three types of criterion examples are detailed as follows:
Inconsistency. We employ the personality extraction model introduced in §$\S$3.2.1, and utilize the personality triplet randomly extracted from $\mathbb{D}^{u}$ to substitute a triplet in $h_{cho}$ to formulate $h_{crt}$. For example, $h_{cho}$ “I am a farmer live in a small town” is transformed into $h_{crt}$ “I am a spaceman live in a small town” by replacing <I, job, farmer> with <I, job, spaceman>, which is extracted from $\mathbb{D}^{u}$.
Fabrication. We encode sentences in the dataset, selecting top-$m$ candidates with the highest semantic similarity to $h_{cho}$. A candidate, $h_{crt}$, is randomly chosen ensuring $\text{Ext}(h_{crt})\cap\text{Ext}(\mathbb{D}^{u})=\emptyset$. For example, from the utterance “My hobbies are watching movies and riding bicycles”, we extract triples <I, hobby, watching movies> and <I, hobby, riding bicycles>. As the triples are not involved in $\text{Ext}(D^{u})$, we can adopt this utterance as $h_{crt}$.
@@ 4.1 Datasets
ConvAI2 Dinan et al. (2020) is a high-quality English dataset focused on personalized dialogues. Each dialogue revolves around a specific profile. The dataset is expanded from the classic PersonaChat Zhang et al. (2018) by crowd workers. Cornell Movie-Dialogs Corpus Danescu-Niculescu-Mizil and Lee (2011) contains over $220,000$ dialogues collected from more than $600$ movies with rich meta-data, offering a diverse range of dialogues between $10,000$ pairs of characters. LIGHT Urbanek et al. (2019) is a large-scale crowdsourced fantasy text adventure game research platform. We extract dialoigues of each character to form the dataset used in the experiments.
@@ 4.2 Baselines
Large Language Models have made great progress recently. We select LLaMA-2-7B-Chat and LLaMA-2-13B-Chat Touvron et al. (2023) as the backbones of IDL, and name the models LLaMA-2-7B IDL and LLaMA-2-13B IDL, respectively. Besides, Vicuna [^4] https://lmsys.org/blog/2023-03-30-vicuna/ and WizardLM Xu et al. (2023) are involved in comparison, where the former is an open-source chatbot developed by fine-tuning LLaMA with ShareGPT, and the latter is fine-tuned from LLaMA-2, starting with a basic set of instructions.
In the ConvAI2 dataset, we compare our IDL models with both profile-based and profile-free approaches. Unlike existing profile-based methods that don’t use Large Language Models (LLMs), we fine-tune LLaMA-2 models (7B and 13B versions) with ConvAI2’s profiles for a fair comparison, naming them LLaMA-2-7B gold and LLaMA-2-13B gold. We also include two other LLM baselines: LLaMA-2 System, which uses profiles directly in system instructions without further training, and LLaMA-2 FT, which fine-tunes on ConvAI2 treating each conversation as a separate example.
For the Movie and LIGHT datasets, we test the adaptability of our IDL models (LLaMA-2-7B IDL and LLaMA-2-13B IDL, both fine-tuned on ConvAI2) against other LLMs using ICL.